How do you find a fraction between a pair of fractions. ex: 7/12, 3/4
Thank you!

Answer by stanbon(48546)

How do you find a fraction between a pair of fractions. ex: 7/12, 3/4
----------------
Express both fractions with the same denominator:
7/12 = 7/12
3/4 = 9/12
--------------
A fraction between 7/12 and 9/12 would be 8/12
=================
Cheers,
Stan H.